OpenStudy (anonymous):

solve for x:

OpenStudy (anonymous):

x = 4

OpenStudy (anonymous):

\[2 \sin ^2 x - coz x -1 = 0 for 0 \le x < 2\pi. \]

OpenStudy (anonymous):

gimmick is to rewrite \[\sin^2(x)=1-\cos^2(x)\] and then get a quadratic in cosine

OpenStudy (anonymous):

\[2(1-\cos^2(x))-\cos(x)-1=0\] \[2-2\cos^2(x)-\cos(x)-1=0\] \[2\cos^2(x)+\cos(x)-1=0\] and then solve as if it was \[2z^2+z-1=0\]

OpenStudy (anonymous):

thanks again for your input satellite73! the question states that the answer should be in terms of pi. does this change the calculation?

OpenStudy (anonymous):

no. just solve for z, and substitute Cos(x) in when you get your answer. Take inverse cosine and you will get your X term in terms of pi.

OpenStudy (anonymous):

thanks mathandphysics! you guys are life-savers! much appreciated.